# Many Heads Are Better Than One: Improved Scientific Idea Generation by A LLM-Based Multi-Agent System

### v2

@@ 5.4 Effects of Different Exploration Mechanisms on Scientific Collaboration
5.5 Effects of Different Underlying LLMs
@@
6 Conclusion
7 Discussions
While our system effectively models scientific collaboration, the simulated interactions may oversimplify real-world teamwork complexities. In large research communities, multiple teams often collaborate dependently or independently on related projects, with agents participating across different teams simultaneously.
To address these limitations, future work could scale up VirSci to a societal scale. Enhancing teamwork simulation by enabling concurrent team activities and allowing agents to contribute to multiple teams or projects would better capture the collaborative dynamics of modern scientific research. These improvements would also provide a more powerful tool for the Science of Science community, enabling deeper exploration of how collaboration drives innovation through dynamic and interactive team processes.
Acknowledgements
Appendix A Effect of the Potential Data Leakage
Appendix B Experimental Settings
B.1 Implementation
We implement our system on top of the Agentscope framework (agentscope, ), which serves for LLM-empowered multi-agent applications. We evaluate our system using different publicly available LLMs: GPT-4o (openai, ), LLaMA3.1-8b (llama3, ), and LLaMA3.1-70b. GPT-4o is accessible exclusively via a public API, while the LLaMA3.1 models are open-weight and invoked using the Ollama (ollama_github, ) in our experiments. LLaMA3.1-8b is chosen as the default LLM considering both efficiency and capability, where more experiments using different LLMs are shown in Sec. 5.5. Each experimental run on LLaMA3.1-8b takes approximately 10 minutes on 1 NVIDIA A100 40G GPU within a team discussion setting of 4 members and 5 turns ($K$ = 5). All experimental results are averaged on 20 runs.
B.2 Dataset Details
B.2.1 Computer Science Dataset.
We first build our scientific research ecosystem using real scientists’ information from the AMiner Computer Science dataset  [^7] https://www.aminer.cn/aminernetwork, which was constructed by extracting scientists’ profiles from online web databases (tang2008arnetminer, ). This dataset includes 1,712,433 authors and 2,092,356 papers, covering the period from 1948 to 2014, with disambiguated author names. To manage the large volume of data, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2000, 2010, and 2014, respectively. For quality assurance, we filtered out past papers lacking abstracts or with fewer than 10 citations, contemporary papers with fewer than 5 citations or missing abstracts, and authors with fewer than 50 papers or 50 co-authors. As a result, we extracted detailed information from 156 authors and 178,197 papers to construct the ecosystem (85,217 papers for the past database and 92,980 papers for the contemporary database) and initialize the corresponding agents for the simulation. All paper and author data are embedded using the “mxbai-embed-large” model (emb2024mxbai, ).
B.2.2 Cross-domain Dataset.
To demonstrate the generalizability and robustness of the system, we conduct additional experiments on the Open Academic Graph 3.1  [^8] https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, which developed from the Open Academic Graph (zhang2022oag, ). This dataset includes 35,774,510 authors and 130,710,733 papers as of 2023, spanning diverse domains such as physics, chemistry, computer science, and biology. To manage this extensive dataset, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2010, 2020, and 2023, respectively. For quality control, we removed past papers without abstracts or with fewer than 200 citations, contemporary papers with fewer than 50 citations or missing abstracts, and authors with fewer than 40 papers or 50 co-authors. This filtering resulted in detailed information on 3,169 authors and 201,131 papers to construct the ecosystem (139,646 papers for the past database and 61,485 papers for the contemporary database) and initialize agents for simulation. Unlike the Computer Science dataset, the Open Academic Graph lacks author-level information on research interests, citations, and co-authors. To address this, we extracted each author’s published papers between $y_{start}$ and $y_{bound}$ and enriched their profiles with the keywords, citations, and authors associated with these papers. Given that the complete set of keywords from all an author’s published papers could be overwhelming, we prompted GPT-4o (openai, ) to extract and summarize these keywords into concise research interests for each author. All paper and author data were similarly embedded using the “mxbai-embed-large” model.
B.3 Detailed Comparison Settings
Given the problem formulation of the AI Scientist, HypoGen, and ours, we must make several justifications to ensure relative fairness in our comparisons: (1) Since the AI Scientist is limited to generating ideas from predefined topics (2D Diffusion, NanoGPT, and Grokking), we include NanoGPT in the topic selection prompt for both HypoGen and VirSci as the initial discussion topic, ensuring that the final abstracts align with the same research direction. (2) Given the different approaches to idea generation, we ensure that comparisons are made under similar inference costs. The AI Scientist performs 50 turns of self-reflection during its idea generation, which does not apply to its paper (abstract) generation. To match the inference costs, we set VirSci with 4 team members and 5 discussion turns, and set HypoGen to 12 discussion turns, ensuring that the experiments are conducted under comparable computational costs. (3) Since the AI Scientist and HypoGen lack a scientific research ecosystem, they retrieve papers across all time ranges through the Semantic Scholar API (fricke2018semantic, ) or PubMed [^9] https://pubmed.ncbi.nlm.nih.gov/. To maintain consistency, we replace our databases and HypoGen’s PubMed API with the Semantic Scholar API for paper retrieval and metric calculation. Specifically, after generating ideas and corresponding abstracts, we use the generated ideas as queries to retrieve related papers, extracting the relevant abstracts and citation counts for evaluation. (4) We evaluate the generated abstracts using both our metrics (CD and CI), the AI Scientist’s metric (LLM review score) (lu2024ai, ) and the human evaluation (si2024can, ). The LLM review score is calculated by GPT-4o, which conducts abstract reviews based on a truncated version of the Neural Information Processing Systems (NeurIPS) conference review guidelines [^10] https://neurips.cc/Conferences/2024/ReviewerGuidelines, shown in Fig. 25.
B.4 Human Evaluation
To comprehensively compare our method with baseline models, we additionally conduct a human evaluation to assess novelty, feasibility, and effectiveness besides our proposed metrics and LLM review score. Ten PhD students in computer science (the relevant field for the selected research topic), unaware of the method identities, rated the abstracts on a 10-point scale (1: Poor, 10: best) based on three metrics (si2024can, ): 1. Novelty: Whether the idea is creative and different from existing works on the topic, and brings fresh insights; 2. Feasibility: How feasible it is to implement and execute this idea as a research project; 3. Effectiveness: How likely the proposed idea is going to work well (e.g., better than existing baselines). The 10-point scale guideline is provided for human experts and is designed based on the principles outlined by si2024can .
Appendix C More Analysis of Proposed Metrics
@@
In this paper, we propose objective evaluation metrics to assess the novelty of system outputs, primarily based on vector similarity. This approach aligns with the research in the Science of Science domain, where our computational methods (relying on vector union and overlap) draw on established literature, such as  boyack2005mapping ; shi2023surprising ; liu2023data .
@@
Another potential concern is the introduction of bias due to a lack of diversity in the datasets or the generated ideas. To address this, we employ two large-scale datasets (Computer Science dataset and Open Academic Graph dataset) in our experiments to ensure content diversity. Additionally, each scenario in our experiments is tested 20 times to ensure that the conclusions regarding the generated ideas are not influenced by potential biases.
@@
To further assess the validity of our proposed overall novelty metric, we selected 200 abstracts generated under various experimental conditions from the Computer Science dataset. These abstracts were evaluated using three approaches: (1) our proposed overall novelty metric, (2) LLM-based reviewers (utilizing the GPT-4o API), and (3) human researchers specializing in computer science (detailed in Appx. B.4). For both LLM-based reviewers and human researchers, we employed the novelty scoring framework from (si2024can, ) as a guideline. To enhance objectivity, each abstract was evaluated three times by LLM-based reviewers, with the average score calculated. Similarly, three independent human researchers assessed each abstract, and their average score was computed. The evaluation results are presented in Fig. 10 and Fig. 11. In Fig. 10, the axes represent the scores of the same abstract evaluated under our proposed metric and the LLM-based reviewers. In Fig. 11, the axes represent the scores of the same abstract evaluated under our proposed metric and by human researchers. The Pearson correlation coefficients between our proposed overall novelty metric and both LLM-based reviewers and human researchers demonstrate a positive correlation with established novelty measurement methods (lu2024ai, ; si2024can, ), which, to some extent, supports the validity of our metric.
Appendix D Discussions on System Feasibility
@@
Figure 13: The recently published similar paper (kuhnisch2022caries, ) of abstract, corresponding to the first example abstract generated by our VirSci. This abstract discusses the application of artificial intelligence in caries management.
@@
Figure 15: The recently published similar paper (cong2022magnetic, ) of abstract, corresponding to the second example abstract generated by our VirSci. This abstract discusses the application of robots in cancer treatment.
To better show the feasibility of the proposed VirSci, we present two example abstracts generated by our system using the Open Academic Graph 3.1 dataset from 2011 to 2020, alongside corresponding similar recently published papers, to highlight the practical relevance and applicability of our system’s outputs. The first example abstract generated by our system is shown in Fig. 12, while the corresponding similar paper (kuhnisch2022caries, ) of the abstract is shown in Fig. 13. This pair of abstracts discusses the application of artificial intelligence in caries management. The second example abstract generated by our system is shown in Fig. 14, while the corresponding similar paper (cong2022magnetic, ) of the abstract is shown in Fig. 15. This pair of abstracts discusses the application of robots in cancer treatment. Overall, these examples demonstrate that our system has the potential to discover novel scientific ideas.
Appendix E Prompts
E.1 Scientist Definition
@@ E.2 Collaboration Selection
E.3 Topic Discussion
E.3.1 Discussion
@@ E.5 Novelty Assessment
E.6 Abstract Generation
E.6.1 Discussion
@@ E.7 LLM Review
Figure 25: LLM review. To ensure a fair comparison, we use the same metric as AI Scientist, which is based on NeurIPS2024 reviewer guidelines. We keep several critical metrics in this guideline since we only evaluate the abstract.
Appendix F Example Scenarios
F.1 Collaboration Selection
@@
F.2 Topic Discussion
F.2.1 Topic Discussion Normal Case
The example scenario of the normal case in the topic discussion is illustrated in Fig. 27. Scientists will provide topic discussion responses, which ensure a coherent discussion of the research topic.
Figure 27: The example scenario of the normal case in the topic discussion. Scientists will provide topic discussion responses (highlighted in yellow), which ensure a coherent discussion of the research topic.
F.2.2 Invitation Mechanism
@@ F.4 Novelty Assessment
F.5 Abstract Generation
F.5.1 Abstract Generation Normal Case

### v3

@@ D.4 Effects of Different Exploration Mechanisms on Scientific Collaboration
D.5 Effects of Different Underlying LLMs
@@
Appendix E More Analysis of Proposed Metrics
@@
In this paper, we propose objective evaluation metrics to assess the novelty of system outputs, primarily based on vector similarity. This approach aligns with the research in the Science of Science domain, where our computational methods (relying on vector union and overlap) draw on established literature, such as  Boyack et al. (2005); Shi and Evans (2023); Liu et al. (2023).
@@
Another potential concern is the introduction of bias due to a lack of diversity in the datasets or the generated ideas. To address this, we employ two large-scale datasets in our experiments to ensure content diversity. Additionally, each scenario in our experiments is tested 20 times to ensure that the conclusions regarding the generated ideas are not influenced by potential biases.
To further assess the validity of our proposed overall novelty metric, we selected 200 abstracts generated under various experimental conditions from the Computer Science dataset. These abstracts were evaluated using three approaches: (1) our proposed overall novelty metric, (2) LLM-based reviewers (utilizing the GPT-4o API), and (3) human researchers specializing in computer science (detailed in Appx. C.4). For both LLM-based reviewers and human researchers, we employed the novelty scoring framework from (Si et al., 2024) as a guideline. To enhance objectivity, each abstract was evaluated three times by LLM-based reviewers, with the average score calculated. Similarly, three independent human researchers assessed each abstract, and their average score was computed. The evaluation results are presented in Fig. 10 and Fig. 11. In Fig. 10, the axes represent the scores of the same abstract evaluated under our proposed metric and the LLM-based reviewers. In Fig. 11, the axes represent the scores of the same abstract evaluated under our proposed metric and by human researchers. The Pearson correlation coefficients between our proposed overall novelty metric and both LLM-based reviewers and human researchers demonstrate a positive correlation with established novelty measurement methods (Lu et al., 2024; Si et al., 2024), which, to some extent, supports the validity of our metric.
@@
Appendix F Discussions on System Feasibility
To better show the feasibility of the proposed VirSci, we present two example abstracts generated by our system using the Open Academic Graph 3.1 dataset from 2011 to 2020, alongside corresponding similar recently published papers, to highlight the practical relevance and applicability of our system’s outputs. The first example abstract generated by our system is shown in Fig. 12, while the corresponding similar paper (Kühnisch et al., 2022) of the abstract is shown in Fig. 13. This pair of abstracts discusses the application of artificial intelligence in caries management. The second example abstract generated by our system is shown in Fig. 14, while the corresponding similar paper (Cong et al., 2022) of the abstract is shown in Fig. 15. This pair of abstracts discusses the application of robots in cancer treatment. Overall, these examples demonstrate that our system has the potential to discover novel scientific ideas.
Appendix G Consistency Between Two Datasets.
As shown in Fig. 4, no significant differences were found in outputs from single-agent teams (team size equals 1) across the two datasets, as individual research lacks interdisciplinary input. The overall trend—novelty rising, peaking at a team size of 8 and 5 turns, then declining—was consistent across both datasets, demonstrating our platform’s robustness and the potential of multi-agent systems in enhancing scientific idea generation. These findings also support the value of interdisciplinary collaborations for driving higher-impact research (Shi and Evans, 2023).
Figure 13: The recently published similar paper (Kühnisch et al., 2022) of abstract, corresponding to the first example abstract generated by our VirSci. This abstract discusses the application of artificial intelligence in caries management.
@@ Appendix F Discussions on System Feasibility
Figure 15: The recently published similar paper (Cong et al., 2022) of abstract, corresponding to the second example abstract generated by our VirSci. This abstract discusses the application of robots in cancer treatment.
Appendix H Prompts
H.1 Scientist Definition
@@ H.2 Collaboration Selection
H.3 Topic Discussion
H.3.1 Discussion
@@ H.5 Novelty Assessment
H.6 Abstract Generation
H.6.1 Discussion
@@ H.7 LLM Review
Figure 25: LLM review. To ensure a fair comparison, we use the same metric as AI Scientist, which is based on NeurIPS2024 reviewer guidelines. We only keep several critical metrics in this guideline since now we only need to evaluate the abstract.
Appendix I Example Scenarios
I.1 Collaboration Selection
@@
I.2 Topic Discussion
I.2.1 Topic Discussion Normal Case
The example scenario of the normal case in the topic discussion is illustrated in Fig. 27. Scientists contribute responses based on the discussion history and their individual knowledge, leading to distinct agent behaviors that reflect their diverse expertise profiles.
Figure 27: The example scenario of the normal case in the topic discussion. Scientists generate responses based on the discussion history and their own knowledge (highlighted in yellow), promoting a diverse and informed exploration of the research topic.
I.2.2 Invitation Mechanism
@@ I.4 Novelty Assessment
I.5 Abstract Generation
I.5.1 Abstract Generation Normal Case